please help me solve this equation: 15y2 + 45y - 9 = 4y - 5y2
Answer by checkley77(12844)

15y2 + 45y - 9 = 4y - 5y2
15Y^2+5Y^2+45Y-4Y-9=0
20Y^2+41Y-9=0
(5Y-1)(4Y+9)=0
5Y-1=0
5Y=1
Y=1/5
Y=0.2 ANS.
4Y+9=0
4Y=-9
Y=-9/4
Y=-2.25 ANS.